OpenStudy (anonymous):

how do I solve tan 2pi/3

OpenStudy (imstuck):

the very first thing I ALWAYS do when working with these is to convert it to degrees. I don't know about you, but my mind does not naturally work in radians. Degrees I can do. 2pi/3 is converted as such:\[\frac{ 2\pi }{ 3 }\times \frac{ 180 }{ \pi }=120\]So it's 120 degrees. Let's draw that.

OpenStudy (anonymous):

Try looking it up on Wolfram Alpha. They give detailed explanations for mostly every problem in many subjects :)

OpenStudy (anonymous):

@IMStuck okay but when i try solving for tan 5pi/6 I keep getting 1/-sqrt3

OpenStudy (anonymous):

and I know that's wrong because it isn't an answer choice